OpenStudy (anonymous):

How to find the gradient of this line: y = 4 – 3x

OpenStudy (anonymous):

There exists 2 famous methods one is algebraic and the other is calculus, regardless of other methods such as the graphical method

OpenStudy (anonymous):

y = 4 – 3x // i'm pretty sure that by gradient, they mean what is the slope of the line // rewrite in the form y = mx+b, where m is the slope/gradient. y = (-3)x+4 your slope/gradient is -3. attached is a graph of your line http://www.wolframalpha.com/input/?i=y+%3D+-3x%2B4

OpenStudy (anonymous):

you may differentiate the linear equation you have

OpenStudy (anonymous):

y = 4-3x, -> dy/dx= -3=m

OpenStudy (anonymous):

or use the superposition principle to compare the above equation with the general linear form of a linear equation, which is y = mx+c

OpenStudy (anonymous):

y = -3x+4, therefore m=-3

OpenStudy (anonymous):

there exist a lot of methods to solve this problem